P(x) = (u^x. e^-u) / x! and let u=10. Find P(9).


1
Expert's answer
2021-03-02T08:11:17-0500
"P(x) = \dfrac{u^x e^{-u}}{x!}\\"

at u = 10 and x =9



"P(9) = \dfrac{10^{9} \times e^{-10}}{9!}\\
= \dfrac{10^{9} }{9! \times e^{10}}\\
= \dfrac{10^{9} }{362880 \times 22026.47}\\
= \dfrac{10^{9} }{7992965433.6}
= \dfrac{1000000000 }{7992965433.6}\\
\simeq 0.12511 \\\implies 125.11 \times 10^{-3}"